OpenStudy (anonymous):

translate the following into an equation, do NOT slove.... Cheryl's commute is 4.5 miles further than Joanna's. Matthews commute is 21 miles less than 5 times times Joanna's commute. If their average commute is 13 miles, how far is Joanna's commute? Is my answer even close??? Joanna's commute will be x x + x + 4.5 + x * 5 - 21 divided by 3

OpenStudy (amistre64):

c = j+4.5 m = 5j-21 average commute = 13.... (m+j+c)/3 (j+4.5) + (5j-21) + j ------------------ 3

OpenStudy (amistre64):

why use x for joanna? j is so much easier to relate to in this

OpenStudy (amistre64):

(j+4.5) + (5j-21) + j ------------------ = 13 ;) 3